Circular No. 6113

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPENOVA 1994ag IN ANONYMOUS GALAXY R. H. McNaught reports his discovery of an apparent supernova (mag

∼ 17.5) in an anonymous galaxy on an I plate taken by M. Hartley with the U.K. Schmidt Telescope on Nov. 11. SN 1994ag is at α = 3^{h}51^{m}42^{s}.88, δ = −39^{o}09′16′′.6 (equinox 1950.0; uncertainty 0′′.3 in each coordinate), which is 8′′.6 east and 3′′.1 south of the galaxy’s center. No image appears in this position on the ESO B survey. The galaxy has a pronounced nucleus with a bar, the supernova appearing at one end of the bar. Confirmation was obtained on Dec. 5 by G. J. Garradd (Loomberah, N.S.W.) via a blue CCD image exposed with a 0.25-m reflector. A nearby star some two mag fainter in B than SN 1994ag has position end figures 44^{s}.08, 20′′.8.

COMET MACHHOLZ (1994r)

Improved parabolic orbital elements from MPC 24368:

T = 1994 Oct. 2.5858 TT  ω = 142.^{o}7985
Ω = 249.9431 }2000.0
q = 1.845571 AU  i = 101.7434
1994/95 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 14 3^{h}48.^{m}96 +19^{o}41.^{′}1 1.113 2.057 157.^{o}5 10.^{o}5 10.9
19 3 34.68 +15 00.3 1.185 2.085 147.8 14.6 11.1
24 3 23.34 +10 55.8 1.273 2.114 138.7 17.9 11.3
29 3 14.49 + 7 28.1 1.372 2.144 130.4 20.5 11.5
Jan. 3 3 07.74 + 4 34.2 1.481 2.176 122.7 22.4 11.7
8 3 02.77 + 2 09.8 1.597 2.209 115.7 23.7 12.0
13 2 59.26 + 0 10.3 1.717 2.242 109.2 24.5 12.2
18 2 56.97 − 1 28.5 1.841 2.277 103.2 24.9 12.4
23 2 55.71 − 2 50.4 1.967 2.312  97.6 25.0 12.6
28 2 55.31 − 3 58.5 2.093 2.349  92.3 24.8 12.8
Feb. 2 2 55.66 − 4 55.4 2.219 2.386  87.3 24.4 13.0
7 2 56.64 − 5 43.1 2.344 2.424  82.6 23.8 13.2
12 2 58.15 − 6 23.5 2.467 2.462  78.1 23.1 13.4
17 3 00.13 − 6 58.1 2.589 2.501  73.9 22.3 13.5
22 3 02.51 − 7 27.9 2.707 2.541  69.8 21.4 13.7

1994 December 12 Daniel W. E. Green